# A projection method for particle resampling

### v6

@@ Abstract
Particle discretizations of partial differential equations are advantageous for high-dimensional kinetic models in phase space due to their better scalability than continuum approaches with respect to dimension. Complex processes collectively referred to as particle noise hamper long time simulations with particle methods. One approach to address this problem is particle mesh adaptivity or remapping, known as particle resampling. This paper introduces a resampling method that projects particles to and from a (finite element) function space. The method is simple; using standard sparse linear algebra and finite element techniques, it can adapt to almost any set of new particle locations and preserves all moments up to the order of polynomial represented exactly by the continuum function space.
This work is motivated by the Vlasov-Maxwell-Landau model of magnetized plasmas with up to six dimensions, $3X$ in physical space and $3V$ in velocity space, and is developed in the context of a $1X$ + $1V$ Vlasov-Poisson model of Landau damping with logically regular particle and continuum phase space grids. Stable long time dynamics are demonstrated up to $T=500$ and reproducibility artifacts and data with stable dynamics up to $T=1000$ are publicly available.
@@ 1 Background
Particle, marker particle, or macro-particle methods, such as particle-in-cell (PIC), are discretizations that, akin to traditional continuum-based methods such as finite elements (FE), finite volume, etc., discretize continuous PDE models as opposed to discrete, ground truth models like molecular dynamics. Particle methods scale with dimension with a theoretical accuracy of $\mathcal{O}(N^{-\frac{1}{2}})$ and complexity $\mathcal{O}(N)$, with $N$ particles, whereas continuum methods have higher order accuracy, $\mathcal{O}(N^{-p})$ for some order $p$, usually two or higher, the complexity is order $\mathcal{O}(N^{D})$ with $N$ grid points in each dimension $D$. For example, with a commonly attainable $p=2$ the complexities cross over at $D=4$ and particle methods have lower order complexity at higher dimensions. This scaling with dimension has motivated the use of the PIC methods for the Vlasov-Maxwell-Landau (VML) system, or Boltzmann’s equation with Coulomb collisions in Landau form for magnetized plasmas and Vlasov-Poisson-Landau in astrophysics [4].
@@
Mesh or grid adaptivity is a fundamental tool in PDE modeling for both continuum and particle grid methods. Particle adaptivity, known as particle resampling in the physics community. This paper develops a particle resampling approach that uses a conservative mapping between particles and continuum grids, a projection [36], that conserves an arbitrary number of moments exactly, that was developed by the structure preserving discretization community for analysis purposes [22, 44]. This method allows for remapping a particle distribution to any new set of particles, supporting adaptivity in both the continuum grid construction and the particle grid definition. The focus of this paper is investigating the use of this projection for particle resampling with static regular grids for both particles and the finite element space with a standard plasma model problem. The testing codes are built on PETSc (Portable, Extensible Toolkit for Scientific Computation), and are publicly available (Appendix §A).
@@
Particle resampling methods have been developed by many groups. Lapenta developed a method in the 1990s with a solve of the form, $\left(M_{p}M_{p}^{T}\right)^{-1}$, with Lagrange multipliers that enforce moment constraints explicitly [26, 25, 27], that is formally similar to the pseudoinverse solve in our approach §2.2. Colella et al., developed a particle remapping method that is similar to our approach in a finite volume context with a direct remap from the grid back to particles [43, 32, 33]. Faghihi et al. developed resampling methods with moment constraints and linear programming to enforce moments and other algebraic constraints [12]. Gonoskov proposed probabilistic down-sampling algorithms using algebraic constraints to enforced conservation [15]. Pfeiffer et al. introduced two conservative particle split and merge methods that use statistical properties of the plasma such as thermal speed [34]. Several groups have presented particle coalescence and splitting schemes, often using trees, on small groups of locally binned particles [39, 42, 5, 45, 28].
@@
Our approach is distinguished from previous work in a few ways. An arbitrary number of moments are implicitly and provably conserved exactly without resorting to explicitly constraining the solution to desired manifolds. Our method is simple and built entirely on standard sparse linear algebra primitives and allows for almost any new distribution of particles, the original particle grid being a good choice if dynamic adaptivity is not desired. The Colella et al. algorithm is also flexible with phase-space adaptive mesh refinement (AMR) and is similar to our algorithm if the resampling set is chosen to align with the vertex points of the grid ($M_{p}=I$ in §2.1) [33], which results in the key component of this projection method, the pseudoinverse, vanishing.
@@
This paper proceeds with relevant background in structure preserving methods in §2, the projection based resampling method in §3, numerical methods and the test problem in §4, §5 experiments with a direct remap method. Experiments with the full high-order finite element projection and remapping method are presented in §6, §7 investigates long time behavior with convergence studies, §8 discusses side effects of resampling, and §9 concludes with a discussion of potential future work.
@@ 2 Structure preserving methods for Boltzmann’s equations
The critical idea in this work comes from research on structure preserving methods for Boltzmann’s equations in general and the VML system for magnetized plasmas in particular, that results in a simple and elegant algorithm that provably conserves an arbitrary number of moments.
Hamiltonian models in phase space where density is a function of both space ($\mathbf{x}$) and velocity or momentum space ($\mathbf{v}$) are the fundamental equation of gravitational dynamics and electrostatics plasmas with the Vlasov-Poisson system, and electromagnetic plasmas with the Vlasov-Maxwell system. A Coulomb collision term accounts for the statistics of particle interactions not present in the Hamiltonian [16, 18, 3, 41, 20], giving rise to the governing equations for magnetized plasmas where the density of each species $\alpha$ is evolved in phase space according to
@@
where the collisional term is summed over all species $\beta$. This equation is composed of the symplectic Vlasov-Maxwell term $\frac{df}{dt}=0$ and a metric, or diffusive, collision operator $C$, and Maxwells’s equations provide an expression for $\frac{\partial\vec{v}}{\partial t}=a=\frac{q_{\alpha}}{m_{\alpha}}\left(\mathbf{E}+\mathbf{v}\times\mathbf{B}\right)$.
This system has rich mathematical structure that can be preserved with proper discretizations. The metriplectic formalism is an approach to analyze VML and to develop structure preserving discretizations for the VML system [18, 22, 17]. When a structure preserving grid-based collision operator [18, 2, 1, 3] is coupled with a PIC method, a mechanism is needed to map distribution functions, in velocity space, between a particle representation and a finite element basis representation of the distribution function that preserves moments, as well as other structure [36]. Preserving the second moment in velocity space, energy, is critical for many applications.
@@ 2.1 Structure preserving particle-finite element basis mapping
To apply a continuum operator in a PIC method that conserves moments a conservative particle-finite element basis mapping and remapping method is required. Given a particle with weight $w_{p}$ and position $\textbf{x}_{p}$, a delta function representation $f_{p}(\textbf{x})=w_{p}\delta(\textbf{x}-\textbf{x}_{p})$, and a finite element (FE) space $V$ of functions $\phi_{i}$ and coefficients $\rho_{i}$, a function can be expressed as $f_{FE}(\textbf{x})=\sum_{i}\rho_{i}\phi_{i}(\textbf{x})$. Ideally $f_{FE}(\textbf{x})=f_{p}(\textbf{x})$, but that not possible. Weak equivalence can however be enforced with:
[…]
@@
which defines an equation for particle deposition:
@@
This mapping is proven to conserve moments up to the order polynomial that the FE space can represent exactly [19, 36], for example a quadratic element mesh is sufficient to conserve energy in velocity space.
After deposition on the FE space, a Poisson or Ampere’s law solve can be executed or a collision operator, $L$, can be evolved, $u\leftarrow L\rho$. In mapping $u$ back to particles one can simply invert (2), $\overline{w}\leftarrow M_{p}^{-1}Mu$, however $M_{p}$ is rectangular in general. The key idea is that a pseudoinverse, $M_{p}^{\dagger}$ with $M_{p}M_{p}^{\dagger}=I$, conserves moments:
@@
Thus, if the collision operator conserves moments this entire process of applying a continuum operator in a PIC method conserves moments [18].
@@ 2.2 Pseudoinverses and idempotent projections
There are two basic approaches to the pseudoinverse: an appropriate Krylov methods such as LSQR or Moore-Penrose. Both of these solvers are $l_{2}$ projections, but there are alternative norms such as $L_{2}$ that could be investigated. Moore-Penrose is attractive because it is easier to precondition a square matrix, especially for batch solvers [3]. Krylov methods are attractive as they can solve singular systems transparently. Preconditioning LSQR requires some effort, but the pseudoinverse solves in this work are very well conditioned and unpreconditioned LSQR works well. Though we use LSQR in this work, the analysis is clearer with the Moore-Penrose pseudoinverse §3.1 because it is defined with standard matrix inverse: $M_{p}^{\dagger}\equiv M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}$, and stability is easier to understand.
@@
which is a type of “coarse-graining” algorithm, a mechanism to add numerical entropy dissipation [8, 40].
@@
Information is lost while projecting a particle representation of a function onto a FE basis if the number of particles exceeds the number of FE basis functions, which is the case of interest. An attractive property of (4) is that information is only lost on the first application in that
@@ 3 A particle resampling method
The key observation is that after computing (2), the distribution function is entirely represented on the FE space and particle weights and positions are no longer needed. A new set of particle positions, essentially any new set, can be created. A new particle mass matrix, $\overline{M}_{p}$, can then be computed and (4) can be continued with $\overline{M}_{p}$. Moments are conserved because it is provable, and experimentally demonstrated, that the projection to the grid preserves moments and the projection from the grid preserves moments. The framework for this resampling method is to rearrange (4), by projecting back to a new set of particles after the deposition according to:
@@ 3.1 Moore-Penrose stability
Care must be taken in the explicit inverse of $M_{p}M_{p}^{T}$ as it can be singular from, for example, empty rows of $M_{p}$ if there are no particles in the support of an FE basis function. Defining the particle grid is under the control of the algorithm, unlike in coarse-graining and field solves, and sufficient constraints must be understood. A necessary condition for stability of the pseudoinverses is that there does not exist a set of vertices whose union of support of associated FE basis functions contains less particles than the number of vertices in the set. This is not rigorous but comes from the intuition that no set of equations (rows of $M_{p}$) should have less non-empty columns than rows, otherwise the matrix locally singular. Further we find that more particles than basis functions are required for stability in, for example, the $1D$ periodic direction if, with degree $Q$ element (e.g., $Q=2$, a Q2 element), there are $Q$ particles per cell, which results in an equal number of particles and basis functions (equations or vertices). $M_{p}$ is square in this case, but we observe that $\left(M_{P}M_{p}^{T}\right)$ is singular except for the special case where the particles and mesh points are aligned as in the direct remap method in §5. This criterion is not practical to check with an arbitrary set of particles, but we believe that a simple criterion is robust in our experience although not tight. With tensor elements, we use $\left(Q+1\right)^{D}$ particles per cell.
4 Numerical methods and Landau damping test
Consider the classic one-dimensional plasma test, Landau damping, given by the initial state,
(5) $f(x,v)=\frac{1}{\sqrt{2\pi}}e^{-v^{2}}\left(1+A\cos\left(kx\right)\right),$
where $k=0.5$, $v_{max}=6$, and we consider three values for the wave amplitude, $A$: $0.5$, $0.01$ and $0.0001$. The Landau damping test is a popular choice for Vlasov benchmarking because it involves a number of purely kinetic effects, such as phase mixing, and it has simple analytical solutions. A detailed description of the analytical solutions to the Landau damping problem can be found in [13]. To focus on the electrostatic kinetic effects, we ignore collisional dynamics and reduce the full VML equations to the collisionless, magnetic-free Vlasov-Poisson (VP) system.
The two cases where the wave amplitude, $A$, is $0.01$ and $0.5$ are referred to as the linear and nonlinear landau damping test cases, respectively [21, 9]. A third test case, $A=0.0001$, sits well within the linear regime and is also common. In the linear case, the small field perturbation is damped out at a rate of $\gamma=-0.153$ in favor of a more homogeneous field. However, in collisionless tests the growth of subgrid modes can disrupt the field damping and cause large gradients to develop in the phase space. These large gradients lead to a sudden regrowth of the field. Previous work [21] has show that the inclusion of collisions can remove these subgrid modes, damping the field to machine precision with a continuum code.
In the nonlinear case, $A=0.5$, the field decay reverses much earlier and the general dynamics differ from that of the linear case. The two primary explanations for this earlier field resurgence are the stronger interaction between the potential well created by the electric field which resonates with and accelerates more of the particles, and the increased phase mixing, evident in phase-space diagrams. To fully capture these dynamics, the nonlinearized form of the Vlasov equation must be considered. Significant work was done by Villani and Mouhot in [31] to analyze the nonlinear Vlasov equation and show that, while the nonlinear dynamics present in this system lead to a weaker initial decay of the electric field, over long enough time scales, the field will damp out, as it does in the linear case. Thus, it is vital to develop the tools necessary to capture the long time evolution of these kinetic plasmas structures. From [21] and [9], we expect the initial damping rate of the field to be $\gamma_{1}=-0.286$ which quickly turns into a field growth at a rate of $\gamma_{2}=0.086$. We will use these values to verify our tests in later sections.
4.1 PETSc test harness

### v9

@@ Abstract
Particle discretizations of partial differential equations are advantageous for high-dimensional kinetic models in phase-space due to their better scalability than continuum approaches with respect to dimension. Complex processes collectively referred to as particle noise hamper long time simulations with particle methods. One approach to address this problem is particle mesh adaptivity, or remapping, known as particle resampling and remeshing.
This work introduces a resampling method that projects particles to and from a (finite element) function space. The method is simple, using standard sparse linear algebra and finite element techniques, and it preserves all moments up to the order of a polynomial represented exactly by the continuum function space. It is distinguished from most other mesh-based methods in that new particle positions and number are decoupled from the mesh, allowing particle and continuum meshes to be adapted relatively independently. While this work is developed with structured particle and continuum phase-space grids on $1X$ + $1V$ Vlasov-Poisson models of Landau damping and two-stream instability, the method is well-suited to unstructured grids. Stable long time dynamics are demonstrated up to time $T=500$. Reproducibility artifacts and data are publicly available.
@@ 1 Background
Particle, marker particle, or macro-particle methods, such as particle-in-cell (PIC), are discretizations that, akin to traditional continuum-based methods such as finite elements (FE), finite volume, etc., discretize continuous PDE models as opposed to discrete, ground truth models, like molecular dynamics. Particle methods scale with dimension with a theoretical accuracy of $\mathcal{O}(N^{-\frac{1}{2}})$ and complexity $\mathcal{O}(N)$, for $N$ particles. By comparison, continuum methods have higher order accuracy, $\mathcal{O}(N^{-p})$ for some order $p$, usually two or higher, and complexity $\mathcal{O}(N^{D})$, with $N$ grid points in each dimension $D$. For example, with a commonly attainable $p=2$ the complexities cross over at $D=4$ and particle methods have lower order complexity at higher dimensions. This scaling with dimension has motivated the use of the PIC methods for the Vlasov-Maxwell-Landau (VML) system, or Boltzmann’s equation with Coulomb collisions in Landau form for magnetized plasmas and Vlasov-Poisson-Landau in astrophysics [Lancellotti01112007]. Continuum methods are a viable options as well [Gkeyll, Shi2022], and hybrid PIC / continuum methods have been developed [KrausError2017, adams2024performance]. In fact, the method developed here was motivated by conservative coupling of FE and PIC methods and results in a hybrid kinetic, continuum model.
@@
Mesh or grid adaptivity is a fundamental tool in PDE modeling for both continuum and particle methods, and is known as particle resampling in particle methods. This paper develops a particle resampling approach that uses a conservative mapping between particles and continuum grids, a projection [PusztayKnepleyAdams2022], that conserves an arbitrary number of moments exactly, that was developed by the structure preserving discretization community [Kraus2017, WEINSTEIN1981235]. This is flexible, allowing remapping to almost any new set of particles. The focus of this paper is on investigating the efficacy of this method at reducing particle noise, adapting to improve grid quality, as well as undersirable side effects with standard plasma model problems. The testing codes are built on PETSc (Portable, Extensible Toolkit for Scientific Computation), and are publicly available (Appendix §A).
@@
Particle resampling methods have been developed by many groups. Lapenta developed a method in the 1990s with a solve of the form $\left(M_{p}M_{p}^{T}\right)^{-1}$ with Lagrange multipliers that enforce moment constraints explicitly [Lapenta1994213, Lapenta2002317, Lapenta1995139], that is formally similar to the pseudoinverse solve in our approach §2.2. Colella et. al., developed a particle remapping method that is similar to our approach in a finite volume context with a direct remap from the grid back to particles [Colella2011, Myers2017, Myers16]. The Vortex-In-Cell community use resampling, or “remeshing”, extensively, for example, Morgenthal and Walther develop a method similar to Colella et. al. with high-order interpolation to a regular grid that preserves relevant moments and refers to much previous work in remeshing [morgenthal2007a]. Faghihi et al. developed resampling methods with moment constraints and linear programming to enforce moments and other algebraic constraints [FAGHIHI2020109317]. Gonoskov proposed probabilistic down-sampling algorithms using algebraic constraints to enforced conservation [GONOSKOV2022108200]. Pfeiffer et al. introduced two conservative particle split and merge methods that use statistical properties of the plasma such as thermal speed [PFEIFFER20159]. Several groups have presented particle coalescence and splitting schemes, often using trees, within small groups of binned particles [Teunissen2014318, Vranic201565, Assous2003550, Welch2007143, Luu2016165].
@@
Our approach is distinguished from previous works in a few ways. An arbitrary number of moments are implicitly and provably conserved exactly, with exact linear solvers, without resorting to explicitly constraining the solution to desired manifolds. Our method is simple and built entirely on standard sparse linear algebra primitives and allows for almost any new distribution of particles, the original particle grid being a good choice if dynamic adaptivity is not desired. The Colella et. al. algorithm uses phase-space adaptive mesh refinement to provide flexibility in particle positions and is similar to our algorithm if the resampling set is chosen to align with the vertices of the grid (this is explored in §5) [Myers16], which removes the need for a pseudoinverse.
@@
This paper proceeds with relevant background in structure preserving methods in §2, the projection based resampling method in §3, numerical methods and the test problems in §4, and §5 experiments with a direct remap method. Experiments with the full high-order finite element projection and remapping method are presented in §6, §7 investigates long time behavior with convergence studies, §8 discusses side effects of resampling, and §9 concludes with a discussion of potential future work.
@@ 2 Structure preserving methods for Boltzmann’s equations
The critical idea in this work comes from research on structure preserving methods for Boltzmann’s equations in general and the VML system for magnetized plasmas in particular. Hamiltonian models in phase-space where density is a function of both space ($\mathbf{x}$) and velocity or momentum space ($\mathbf{v}$) are the fundamental equation of gravitational dynamics and electrostatics plasmas with the Vlasov-Poisson system, and electromagnetic plasmas with the Vlasov-Maxwell system. A Coulomb collision term accounts for the statistics of particle interactions not present in the Hamiltonian [hazeltine2013plasma, Hirvijoki2017, adams2024performance, Vocks_2002, 2021APSDPPUP1078K], giving rise to the governing equations for magnetized plasmas where the density of each species $\alpha$ is evolved in phase-space according to
@@
where the collisional term is summed over all species $\beta$. This equation is composed of the symplectic Vlasov-Maxwell term $\frac{df}{dt}=0$ and a metric, or diffusive, collision operator $C$. Maxwell’s equations provide an expression for $\frac{\partial\vec{v}}{\partial t}=a=\frac{q_{\alpha}}{m_{\alpha}}\left(\mathbf{E}+\mathbf{v}\times\mathbf{B}\right)$.
This system has rich mathematical structure that can be preserved with proper discretizations. The metriplectic formalism is an approach to analyze VML and to develop structure preserving discretizations [Hirvijoki2017, Kraus2017, Hirvijoki2021]. When a structure preserving grid-based collision operator [Hirvijoki2017, AdamsHirvijokiKnepleyBrownIsaacMills2017, Adams2022a, adams2024performance] is coupled with a PIC method, a mechanism is needed to map distribution functions, in velocity space, between a particle representation and a finite element basis representation that preserves moments, as well as other structure [PusztayKnepleyAdams2022]. Preserving the second moment in velocity space, energy, is critical for many applications.
@@ 2.1 Structure preserving particle-finite element basis mapping
To apply a continuum operator in a PIC method that conserves moments a conservative particle-finite element basis mapping and remapping method is required. Given a particle with weight $w_{p}$ and position $\textbf{x}_{p}$, a delta function representation $f_{p}(\textbf{x})=w_{p}\delta(\textbf{x}-\textbf{x}_{p})$, and a finite element (FE) space $V$ of functions $\phi_{i}$ and coefficients $\rho_{i}$, a function can be expressed as $f_{FE}(\textbf{x})=\sum_{i}\rho_{i}\phi_{i}(\textbf{x})$. Ideally $f_{FE}(\textbf{x})=f_{p}(\textbf{x})$, but that is not possible. Weak equivalence can however be enforced with:
(1) $\int_{\Omega}d{\textbf{x}}~\phi_{j}(\textbf{x})\sum_{p}f_{p}(\textbf{x})=\int_{\Omega}d{\textbf{x}}~\phi_{j}(\textbf{x})\sum_{p}w_{p}\delta(\textbf{x}-\textbf{x}_{p})=\int_{\Omega}d{\textbf{x}}~\phi_{j}(\textbf{x})f_{FE}(\textbf{x})=\int_{\Omega}d{\textbf{x}}~\phi_{j}(\textbf{x})\sum_{i}\rho_{i}\phi_{i}(\textbf{x})\forall\phi_{j}\in V.$
With a particle mass matrix $M_{p}[i,p]\equiv\phi_{i}(\mathbf{x_{p}})$, an FE mass matrix $M[i,j]\equiv\int_{\Omega}d{\textbf{x}}~\phi_{i}(\textbf{x})\phi_{j}(\textbf{x})$, a vector of particle weights $w$ and vector of FE weights $\rho$, (1) can be written in matrix form as
@@
which defines an equation for particle deposition on to the FE space:
@@
This mapping is proven to conserve moments up to the order polynomial that the FE space can represent exactly [KrausError2017, PusztayKnepleyAdams2022], for example a quadratic element mesh in velocity space is sufficient to conserve energy.
After deposition on the FE space, a Poisson or Ampere’s law solve can be applied or a collision operator, $L$, can be evolved, $u\leftarrow L\rho$. In mapping $u$ back to particles one can simply invert (2), $\overline{w}\leftarrow M_{p}^{-1}Mu$, however $M_{p}$ is rectangular in general. The key idea is that a pseudoinverse, $M_{p}^{\dagger}$ with $M_{p}M_{p}^{\dagger}=I$, conserves moments with standard sparse linear algebra:
@@
Thus, if the collision operator conserves moments this entire process of applying a continuum operator in a PIC method conserves moments [Hirvijoki2017].
@@ 2.2 Pseudoinverses and idempotent projections
There are two basic approaches to the pseudoinverse: an appropriate Krylov methods such as LSQR or Moore-Penrose. Both of these solvers are $l_{2}$ projections, but there are alternative norms, such as $L_{2}$, that could be investigated. Moore-Penrose is attractive because it is easier to precondition a square matrix, especially for batch solvers [adams2024performance]. Krylov methods are attractive as they can solve singular systems transparently. Preconditioning LSQR requires some effort, but the pseudoinverse solves in this work are very well conditioned and unpreconditioned LSQR works well. Though we use LSQR in this work, the analysis is clearer with the Moore-Penrose pseudoinverse: $M_{p}^{\dagger}\equiv M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}$, and stability is easier to understand.
@@
which is a type of “coarse-graining” algorithm, a mechanism to add numerical entropy dissipation developed in the physics community [Chen2007-CG, Vernay2012].
@@
Information is lost when projecting a particle representation of a function onto a FE basis if the number of particles exceeds the number of FE basis functions, which is the case of interest. An attractive property of (4) is that information is only lost on the first application of coarse graining in that
@@ 3 A particle resampling method
A key observation is that after computing (2), the distribution function is entirely represented on the FE space and particle weights and positions are no longer needed. A new set of particle positions, essentially any new set, can be created. A new particle mass matrix, $\overline{M}_{p}$, can then be computed and (4) can be continued with $\overline{M}_{p}$. Moments are conserved because it is provable, and experimentally demonstrated, that the projection to the grid preserves moments and the projection from the grid preserves moments. This resampling method rearranges (4) by projecting back to a new set of particles after the deposition according to:
@@ 3.1 Moore-Penrose stability
Care must be taken in the explicit inverse of $M_{p}M_{p}^{T}$ as it can be singular.Stability can be guaranteed because defining the particle grid is under the control of the algorithm, unlike in coarse-graining and field solves. A necessary condition for stability of the pseudoinverses is that there does not exist a set of vertices whose union of support of associated FE basis functions contains less particles than the number of vertices in the set. This is not rigorous but comes from the intuition that no set of equations (rows of $M_{p}$) should have less non-empty columns than rows, otherwise the matrix locally singular. Further we find that more particles than basis functions are required for stability in, for example, the $1D$ periodic direction if, with degree $Q$ element (e.g., $Q=2$, a Q2 element), there are $Q$ particles per cell, which results in an equal number of particles and basis functions (equations or vertices). $M_{p}$ is square in this case, but we observe that $\left(M_{P}M_{p}^{T}\right)$ is singular except for the special case where the particles and mesh points are aligned as in the direct remap method in §5. This criterion is not practical to check with an arbitrary set of particles, and we use $\left(Q+1\right)^{D}$ particles per cell, which is safe and simple, but one could probably create a grid with $Q^{D}+1$ particles in general positions.
3.2 Complexity
The addition of this resampling method in a PIC code results in a hybrid of particle and continuum methods, similar to the effect of using a grid based collision operator in a PIC code [KrausError2017, adams2024performance]. As discussed in §1 with a simple complexity analysis, continuum methods do not scale well with dimension and a more detailed complexity analysis of this method is warranted. In a kinetic plasma model, the resolution requirements in velocity space are much smaller than in configuration space. Meshes are application dependent, but velocity space distributions do not have the potential for the geometric complexity of configuration space in, say, a fusion device or solar coronal jets. There are no boundaries or material properties (e.g., engineered structures like walls) in velocity space and distributions tend toward smooth near-Maxwellians that can be resolved with about $10^{3}$ mesh points [adams2024performance], however sources such as antennae for RF heating [Yajima_2019], alpha particles from fusion processes [Hu_2025] and runaway electrons [Adams2022a], can result in localized structure that may require, say, $10^{4}-10^{5}$ mesh points. A well resolved tokamak model uses about $10^{8}$ points in configuration space. Thus, one can think of $6D$ kinetic models of magnetized plasmas as scaling roughly like a $4.5D$ model (and more like $4D$ models with a gyrokinetic approximation). Note, this observation indicates why continuum methods are feasible for these problems.
The work and memory complexity of the pseudoinverse is composed of the work of building the particle mass matrix and the pseudoinverse solve for each species. For analysis, assume tensor product Cartesian grids of, in the $6D$ case, two $3D$ grids of the same size for simplicity of analysis with tensor elements, and a constant number of particles per phase-space cell. A particle is interpolated to three vertices in each dimension with $Q2$ elements, which results in $3^{6}=729$ non-zeros per row of $M_{p}$ in $6D$. This is substantial and suggests the use of a matrix-free $M_{p}$ solver, which is sufficient for the unpreconditioned LSQR solver used in this work. The solves in the test problems converge to high accuracy, to conserve moments, in about 50 iterations.
Memory complexity
Stability of the pseudoinverse requires there be more particles than vertices. To ensure this reliably, assume $Q+1$ particles per cell in each dimension ($Q$ is the lower bound for stability). There are $NQ$ vertices in each dimension, asymptotically with $N$ cells, which results in a ratio of number of particles to number of vertices of $\left(\frac{Q+1}{Q}\right)^{6}\approx 11$. The lower bound is one. Particle methods use at least $D+1$ words per particle and one word for each species each mesh vertex for number density. The 787-8 work vectors in LSQR would need to be explicitly stored, resulting in roughly the same memory complexity for both grids and particles, depending on the number of particles per vertex and the number of species, which is assumed to be one herein.
Particle work complexity
The work complexity of applying $M_{p}$ is composed of: 1) $\mathcal{O}(D(Q+1)^{D+1})$ flops per cell for the spectral element Jacobian with sum factorization that is amortized with the $(Q+1)^{D}$ particles per cell, 2) $\mathcal{O}(D(Q+1))$ flops per particle to evaluating shape functions, and 3) $\mathcal{O}((Q+1)^{D})$ flops in applying the interpolant to the source vector for each particle. The application of the interpolants is the dominant cost, which is of the same order as the memory complexity that we avoid with matrix-free. This complexity could probably be improved with optimized algorithms.
Particle data movement complexity
The data movement complexity per cell includes 1) reading $D(Q+1)^{D}$ words of cell closure coordinates to compute the element Jacobian, 2) reading $D$ words of particle coordinate data or, with $(Q+1)^{D}$ particles per cell, $D(Q+1)^{D}$ words per cell, to compute the interpolants, 3) reading $(Q+1)^{D}$ words from the source vector and 4) read/write $(Q+1)^{D}$ words from/to the destination vector per cell. Again, this naive result could no doubt be improved with algorithmic development in the future.
4 Numerical tests and the PETSc test harness